OpenStudy (anonymous):

if i add 15g ofNH3 to 1000g of water,what will the melting and boiling point changes be of the resulting solution? Kb=0.52 degrees C/m & Kf=1.86 degrees C/m.

OpenStudy (anonymous):

ΔTF = KF · b · i ΔTb = Kb · b · i

OpenStudy (anonymous):

i got it thanks!

OpenStudy (anonymous):

oh great, finally somebody who doesnt all to be drawn literally!!! :D

OpenStudy (anonymous):

lol